OpenStudy (anonymous):

Find x intercepts? y=3x^2+11x-4. Do I solve for y then substitute?

OpenStudy (anonymous):

set y=0 then solve

OpenStudy (anonymous):

Y=0 for x intercepts.

OpenStudy (anonymous):

let y = 0

OpenStudy (anonymous):

What does that mean?

OpenStudy (anonymous):

\[\min{y = 3 x^2+11 x-4} = -169/12 at x = -11/6 \]

OpenStudy (anonymous):

am i correct ?

OpenStudy (anonymous):

3x(x+4)(-1(x+4)=3x-1=0=1/3 and x+4=0=-4 coming to: -1/3,-4 is this right?

OpenStudy (anonymous):

or is this only y?

OpenStudy (anonymous):

it also -11/6

OpenStudy (anonymous):

and that the possible answer.

OpenStudy (phi):

I assume you made a typo 3x-1=0=1/3 and x+4=0=-4 coming to: -1/3,-4 is this right? because the answer is +1/3 (which you found, but then typed in -1/3) @142343 x= -11/6 is where the parabola reaches a minimum, but its y-value at x= -11/6 is not zero. The question is asking for the 2 points where the parabola crosses the x-axis